# ChEX: Interactive Localization and Region Description in Chest X-rays

### v1

@@ 0.A.2 Prompt Sets for Report Generation
0.A.3 Learned Regional Bias for Pathology Prompts
0.A.4 Query Type for Region Classification and Explanation
Chest ImaGenome (CIG) | ||
Region Classification (RC) | Region Explanation (RE) | |
Query type | [wAUROC] | [Mic-F1-14] |
Boxes (default task) | 70.460.36 | 53.340.43 |
Textual prompt | 70.760.37 | 54.700.41 |
@@ Appendix 0.C Detailed Benchmark Results
0.C.1 Pathology Detection (OD) Tasks
@@
0.C.2 Sentence Grounding (SG) Task
Sentence Grounding (SG) | ||
MS-CXRM | ||
[mIoU] | [mAP] | |
ChEXVM | 47.521.45 | 44.472.21 |
Zeroshot Contrastive Models | ||
BioVIL[4] (masks) M | 22.751.26 | – |
BioVIL[4] (boxes) M | 28.571.31 | 18.621.37 |
CheXzero[70] (masks) M | 11.940.59 | – |
CheXzero[70] (boxes) M | 15.450.67 | 5.940.64 |
Multitask Generative Models | ||
OmniFM-DR[82] VMO | 46.2 | – |
Supervised Visual Grounding | ||
TransVG[9] M | 48.811.45 | 37.652.61 |
TransVG[9] (BioVIL[4]) M | 52.131.73 | 41.242.73 |
TransVG[9] (CheXzero[70]) M | 53.511.53 | 44.052.63 |
@@ 0.D.2.1 Pathology Tokens
For each positive class , i.e. each class that has target bounding boxes, we iteratively apply Hungarian matching using the matching cost , where matching is applied individually for each . We start with an initial matching. If non-matched predicted boxes remain for that class, we apply matching again with the remaining predicted boxes and all target boxes until all predicted boxes are matched, allowing a target box to be matched with several predicted boxes. We apply the box loss to all pairs of matched predicted/target boxes. Additionally, we apply the focal loss to the box scores with positive targets for boxes matched in the first iteration and negative targets for other boxes and all negative classes. Note that we compute the parameter of the focal loss based on the frequency of positive targets per class in the current batch. Overall, the pathology detection loss is computed as .
@@ 0.D.4.3 Full Report Generation (RG)
For RG, we use pre-defined sets of textual prompts. These prompts may include aspects like pathologies or anatomical regions as studied in Sec. 0.A.2. As pathology prompts, we use the set of 14 findings and pathologies from CheXpert[28]. As anatomy prompts we use a subset of 15 anatomical region names from the regions used in CIG, namely the lower, mid, and upper lung zones, hilar structures, apical zones, as well as the clavicle, each for left and right independently, and additional the mediastinum, the cardiac silhouette, and the abdomen. We encode the prompts with the prompt encoder and then apply the prompt detector to predict bounding boxes and ROI tokens for each prompt. The ROI tokens are then used by the sentence generator to predict textual descriptions for each prompt.
Next, we filter the descriptions to include in the final report. We compute the positive probability for each prompt (and the associated description) by classifying each ROI token similarly as in the region classification task using the CheXpert classes excluding no-finding. If any pathology is positive, we consider the prompt and associated description to be positive. We decided to keep all descriptions of pathology prompts but only positive anatomy descriptions (we refer to Sec. 0.A.2).
@@ 0.E.3 Evaluation of Interactive Capabilities
We evaluated the interactive capabilities as studied in Secs. 5.2 and 5.3 on a small subset of the test sets of CIG and MS-CXR. We use only frontal images that have both CIG and MS-CXR targets and exclude samples where the positive classes present in the MS-CXR targets are not present in the CIG targets, i.e., where CIG and MS-CXR targets are not consistent. This results in a total of 98 images.
@@
For Fig. 6b, we apply the CheXbert classifier on predicted sentences and identify the number of positive predicted pathology classes. We refer to the pathology used in the prompt as queried and to all other pathologies as non-queried.

### v2

@@ Appendix 0.C Detailed Benchmark Results
0.C.1 Sentence Grounding (SG) Task
Sentence Grounding (SG) | ||
MS-CXRM | ||
[mIoU] | [mAP] | |
ChEXVM | 47.521.45 | 44.472.21 |
Zeroshot Contrastive Models | ||
BioVIL[4] (masks) M | 22.751.26 | – |
BioVIL[4] (boxes) M | 28.571.31 | 18.621.37 |
CheXzero[71] (masks) M | 11.940.59 | – |
CheXzero[71] (boxes) M | 15.450.67 | 5.940.64 |
Multitask Generative Models | ||
OmniFM-DR[83] VMO | 46.2 | – |
Supervised Visual Grounding | ||
TransVG[9] M | 48.811.45 | 37.652.61 |
TransVG[9] (BioVIL[4]) M | 52.131.73 | 41.242.73 |
TransVG[9] (CheXzero[71]) M | 53.511.53 | 44.052.63 |
In sentence grounding (SG) (Tab. 7), ChEX demonstrates similar performance to the only generative model OmniFM-DR[83] and the best SupVG model TransVG[9] (with CheXzero[71] backbone), where TransVG shows an advantage on the mIoU metric. Notably, TransVG was trained explicitly on this task. ChEX outperforms TransVG without the CheXzero backbone by mAP and is within 1-std on mIoU. The superiority of ChEX over contrastive models underscores the importance of bounding box supervision for accurate localization.
0.C.2 Pathology Detection (OD) Tasks
@@
In pathology detection (OD) (Tab. 8), ChEX is competitive on 2 of 3 tasks, namely NIH8 and MS-CXR. On VinDr-CXR, where bounding box supervision is available, the SupOD model Faster R-CNN [64], trained on this dataset, performs best, outperforming ChEX by . o-DETR[95] performs slightly worse but still outperforms ChEX, while other DETR-style models are on-par or inferior to ChEX (ChEX outperforms DETR-DC5[6] by ). ChEX doubles the performance of the best WSupOD model Loc-ADPD[51] (with BioVIL backbone). On NIH8, all SupOD models perform very poor (ChEX almost doubles their best performance), while the WSupOD model Loc-ADPD is competitive with ChEX. On MS-CXR, the best SupOD baseline o-DETR (with BioVIL backbone), trained on VinDr-CXR, and the best WSupOD model Loc-ADPD are both within 1-std of ChEX.
@@ 0.C.3 Region Classification (RC) Tasks
On region classification (RC) (Tab. 9) tasks, ChEX clearly outperforms all baselines. The best baseline on MS-CXR is the Faster R-CNN with BioVIL backbone, trained on VinDr-CXR, and is outperformed by , while on CIG the best baseline is the contrastive model BioVIL, which is outperformed by .
@@ 0.C.4 Region Explanation (RE) Tasks
Region explanation (RE) can only be performed by a single generative model, namely RGRG[70]. Other report generation models can only predict the full report but cannot provide descriptions for specified regions. On MS-CXR (Tab. 10), our model ChEX performs similar to RGRG on Mic-F1-14 and METEOR but improves by on Mac-F1-14. On CIG (Tab. 11), ChEX outperforms RGRG by on Mic-F1-14, by on Mac-F1-14, and by on METEOR. This result is particularly noteworthy considering RGRG’s explicit training for this task and highlights the importance of the contrastive losses in ChEX. Contrastive baselines exhibit poor performance on these generative tasks, as they rely on sentence retrieval based on region similarity.
@@ 0.C.5 Full Report Generation (RG) Task
For full report generation (RG) on MIMIC-CXR (Tabs. 12 and 13), ChEX sets a new state-of-the-art on the Ex-F1-14, Mic-F1-5+, and Mac-F1-5+ metrics. Considering the commonly used metric Mic-F1-14, the SOTA model MAIRA-1[26] – having 7B parameters – outperforms ChEX – having only 1B parameters – by only . More notably, ChEX outperforms the more than ten times larger 12B parameter Med-PaLM M by and is only outperformed by by the 80 times larger 84B parameter version. Limitations of ChEX can be observed on the Mac-F1-14 metric, where ChEX is outperformed by the SOTA model Med-PaLM M 84B by but also by smaller models like Prompt-MRG[29] (by ) and RaDialog-RG[57] (by ). ChEX, however, improves over RGRG by , indicating that this is a limitation of training based on anatomical regions and is caused by the poor performance of some rarer pathologies (we refer to the supp. material). On the Ex-F1-14 metric, ChEX outperforms the best baseline Prompt-MRG by . Most larger models do however not report this metric. On the Mic-F1-5+ and Mac-F1-5+ metrics, ChEX outperforms the large SOTA model MAIRA-1 by and , respectively. In language-based metrics such as METEOR, ChEX shows relatively low performance ( lower than the best and lower than the second-best model) due to not being explicitly trained for full report generation but generating reports based on a set of queries. While this approach offers advantages (see Sec. 5.4 of the main paper), it may result in reports differing from reference reports. However, it’s important to note that language-based metrics offer limited insights into the factual correctness of generated reports [49].
In Tab. 14, we provide the CE metric F1 per pathology/finding and compare the results between ChEX and the two baselines RGRG and MAIRA-1.
@@ 0.D.2.1 Pathology Tokens
For each positive class , i.e. each class that has target bounding boxes, we iteratively apply Hungarian matching using the matching cost , where matching is applied individually for each . We start with an initial matching. If non-matched predicted boxes remain for that class, we apply matching again with the remaining predicted boxes and all target boxes until all predicted boxes are matched, allowing a target box to be matched with several predicted boxes. We apply the box loss to all pairs of matched predicted/target boxes. Additionally, we apply the focal loss to the box scores with positive targets for boxes matched in the first iteration and negative targets for other boxes and all negative classes. Note that we compute the parameter of the focal loss based on the frequency of positive targets per class in the current batch. The final pathology detection loss is computed as .
@@ 0.D.4.3 Full Report Generation (RG)
For RG, we use pre-defined sets of textual prompts. These prompts may include aspects like pathologies or anatomical regions as studied in Sec. 0.A.4. As pathology prompts, we use the set of 14 findings and pathologies from CheXpert[28]. As anatomy prompts we use a subset of 15 anatomical region names from the regions used in CIG, namely the lower, mid, and upper lung zones, hilar structures, apical zones, as well as the clavicle, each for left and right independently, and additional the mediastinum, the cardiac silhouette, and the abdomen. We encode the prompts with the prompt encoder and then apply the prompt detector to predict bounding boxes and ROI tokens for each prompt. The ROI tokens are then used by the sentence generator to predict textual descriptions for each prompt.
Next, we filter the descriptions to include in the final report. We compute the positive probability for each prompt (and the associated description) by classifying each ROI token similarly as in the region classification task using the CheXpert classes excluding no-finding. If any pathology is positive, we consider the prompt and associated description to be positive. We decided to keep all descriptions of pathology prompts but only positive anatomy descriptions (we refer to Sec. 0.A.4).
@@ 0.E.3 Evaluation of Interactive Capabilities
We evaluated the interactive capabilities as studied in Secs. 5.3 and 5.2 on a small subset of the test sets of CIG and MS-CXR. We use only frontal images that have both CIG and MS-CXR targets and exclude samples where the positive classes present in the MS-CXR targets are not present in the CIG targets, i.e., where CIG and MS-CXR targets are not consistent. This results in a total of 98 images.
@@
For Fig. 6(b), we apply the CheXbert classifier on predicted sentences and identify the number of positive predicted pathology classes. We refer to the pathology used in the prompt as queried and to all other pathologies as non-queried.